Weighted Average Method, FIFO Method, Physical Flow, Equivalent Units Heap Company manufactures a product that passes...

  1. Weighted Average Method, FIFO Method, Physical Flow, Equivalent Units

    Heap Company manufactures a product that passes through two processes: Fabrication and Assembly. The following information was obtained for the Fabrication Department for September:

    1. All materials are added at the beginning of the process.
    2. Beginning work in process had 80,000 units, 30 percent complete with respect to conversion costs.
    3. Ending work in process had 17,000 units, 25 percent complete with respect to conversion costs.
    4. Started in process, 95,000 units.

    Required:

    1. Prepare a physical flow schedule.

    Heap Company
    Physical Flow Schedule
    Units to account for:
    • Started and completed
    • Units, beginning work in process
    • Units, ending work in process
    • Units completed
    • Units, ending work in process
    • Units started
    • Units transferred out
    Total units to account for
    Units accounted for:
    Units completed and transferred out:
    • Started and completed
    • Units completed
    • Units started
    • Units transferred out
    • From beginning work in process
    • Units completed
    • Units started
    • Units transferred out
    • Units completed
    • Units, ending work in process
    • Units started
    • Units transferred out
    Total units accounted for

    Feedback

    2. Compute equivalent units using the weighted average method.

    Weighted average method: Equivalent Units
    Direct Materials
    Conversion Costs

    3. Compute equivalent units using the FIFO method.

    FIFO method: Equivalent Units
    Direct Materials
    Conversion Costs
Answer #1

1. Prepare a physical flow schedule.

Heap Company
Physical Flow Schedule
Units to account for:

Units, beginning work in process

80000

Units started

95000
Total units to account for 175000
Units accounted for:
Units completed and transferred out:

Started and completed

78000

From beginning work in process

80000

Units, ending work in process

17000
Total units accounted for 175000

Feedback

2. Compute equivalent units using the weighted average method.

Weighted average method: Equivalent Units
Direct Materials 158000+17000 = 175000
Conversion Costs 158000+(17000*25%) = 162250

3. Compute equivalent units using the FIFO method.

FIFO method: Equivalent Units
Direct Materials 95000
Conversion Cost (80000*70%+78000+17000*25%) = 138250